Prove that the group G=[a,b] with the defining set of relations a3=e, b7=e, a-1ba=b8 is a cyclic group of order 3?


1
Expert's answer
2022-01-10T18:41:46-0500

Consider the group "G=\langle a,b\rangle" with the defining set of relations "a^3=e,\ b^7=e,\ a^{-1}ba=b^8." Since "b^7=e," we conclude that the equality "a^{-1}ba=b^8" is equivalent to "a^{-1}ba=b^7b=eb=b," and hence (after left multiplication by "a") to "ba=ab." It follows that the elements "a" and "b" commute. Since "3" and "7" are relatively prime, we conclude that the element "ab" is of order 21, that is "(ab)^{21}=e." Therefore, "G" is a cyclic group of order 21.